Hank Itzek manufactures and sells homemade wine, and he wants to develop a standard cost per gallon. The following are required for production of a 60-gallon batch. 2,400 ounces of grape concentrate at $0.08 per ounce 90 pounds of granulated sugar at $0.57 per pound 102 lemons at $0.82 each 240 yeast tablets at $0.28 each 180 nutrient tablets at $0.11 each 3,200 ounces of water at $0.003 per ounce Hank estimates that 6% of the grape concentrate is wasted, 12% of the sugar is lost, and 21% of the lemons cannot be used Compute the standard cost of the ingredients for one gallon of wine. (Round intermediate calculations and final answer to 2 decimal places, e.g. 1.25.) Standard Cost Per Gallon
Answer #1

Cost of grape concentrate required for 60 gallons = (2,400 ounces*$0.08)/(1-0.06)

= $192/0.94 = $204.26 (Cost Before wastage of 6%)

Cost of granulated sugar for 60 gallons = (90 pounds*$0.57)/(1-0.12)

= $51.3/0.88 = $58.30 (Cost before wastage of 12%)

Cost of lemon for 60 gallons = (102 lemons*$0.82)/(1-0.21)

= $83.64/0.79 = $105.87

Calculation of Standard cost per gallon (Amounts in $)

Ingredients Cost for 60 gallon batch (A) Cost per gallon (A/60 gallons)
Grape Concentrate 204.26 3.41
Granulated Sugar 58.30 0.97
Lemon 105.87 1.76
Yeast Tablets (240*$0.28) 67.20 1.12
Nutrient Tablets (180*0.11) 19.80 0.33
Water (3,200*0.003) 9.60 0.16
Total Standard Cost 465.03 7.75

Therefore standard cost per gallon is $7.75 per gallon.